recent study by the American Accounting Association revealed among 30 graduates in accounting 20 students graduating with a major in accounting select public accounting. Suppose we select a sample of 15 recent graduates.What is the probability that exactly two select public accounting?





1
Expert's answer
2022-05-10T13:33:45-0400

Let "X=" the number of select public accounting:"X\sim Bin(n, p)."

Given "n=15, p=20\/30=2\/3, q=1-p=1\/3."


"P(X=2)=\dbinom{15}{2}(\dfrac{2}{3})^{2}(\dfrac{1}{3})^{15-2}"

"\approx0.0002927"